OpenStudy (anonymous):

Use the compound interest formula to compute the Future value. $400 for 8 years at 6% compounded daily (use n=360). Would my n be 360 or (360)(8)=2880?

OpenStudy (gorv):

compounded daily ?? :o

OpenStudy (anonymous):

Yes

OpenStudy (gorv):

that means 1/365 year

OpenStudy (gorv):

and for 8 year 8/365

OpenStudy (gorv):

sorry that will be 360

OpenStudy (anonymous):

Uhh she told us to use 360 whenever we saw daily. But wouldn't you multiply 8 by it?

OpenStudy (gorv):

that means it is componded 360 times in a year

OpenStudy (gorv):

so rate =6/360 and time=360*8

OpenStudy (anonymous):